40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11814     Entry time: Wed Nov 25 22:59:42 2015
Author: yutaro 
Type: Update 
Category: LSC 
Subject: AS table optics realignment 

I slightly changed the orientation of a few mirrors on AS table that are used to make the AS light get into PDs, in order to confirm that the strange behavior of ASDC (I will report later) is not caused by clipping related to these mirrors or miscentering on PDs.

Then output level of ASDC, AS55, and AS165 could have changed.   

So take care of this possible change when you do something related to them. But the relative change of them would be at most several %, I think.

 

ELOG V3.1.3-